Turner Corporation acquired two inventory items at a lump-sum cost of $50,000. The acquisition included 3,000 units of product LF, and 7,000 units of product 1B. LF normally sells for $15 per unit, and 1B for $5 per unit. If Turner sells 1,000 units of LF, what amount of gross profit should it recognize? a. $1,875 b. $5,625. c. $10,000. d. $11,875.
Answer #1

Total sale value

= LF (3,000*15) + 1B (7,000*5)

= 45,000 + 35,000

= 80,000

LF share = (45,000/80,000)*50,000 = 28,125

Cost of LF per unit = 28,125/3,000 = 9.375

Gross profit = (15 - 9.375)*1000

= 5625

Option B is the answer